You need to prepare 65 mL of a 60% soda/syrup "prescription". To do this, yu can use the 80% soda solution and the syrup solution (0% strength). How many mL of each solution do you need to create this final solution?

Answer by stanbon(75887)

You need to prepare 65 mL of a 60% soda/syrup "prescription". To do this, you can use the 80% soda solution and the syrup solution (0% strength). How many mL of each solution do you need to create this final solution?
------------------
Equation:
soda + soda = soda
----
0.80x + 0(65-x) = 0.60*65
-------
Multiply thru by 100 to get:
80x + 0 = 60*65
x = (3/4)65
x = 48.75 mL (amt. of 80% needed)
65-x = 16.25 (amt. of 0% needed)
======================================
Cheers,
Stan H.